The labor force of a fictional economy, Persepolis, is 500, 000 people. Of these, about 50,000 are frictionally and structurally unemployed, and 50,000 are cyclically unemployed. Another 25,000 are underemployed. Based on this information, Persepolis' natural rate of unemployment is:

  • A.

    10 percent.

  • B.

    15 percent

  • C.

    20 percent

  • D.

    25 percent
Answer #1

Answer: C. 20 percentage

Labor force = 500000

Frictional unemployment = 50000

Cyclical unemployment = 50000

The natural rate of unemployment takes both frictional and cyclical unemployment into account.

So, natural unemployment = 50000 +50000 = 100000

Natural rate of unemployment = Natural unemployment / total labor force

= 100000/500000

= 20 percent